OpenStudy (anonymous):

f(x)=6*x^2-x f(x+h)-f(x)/h

OpenStudy (anonymous):

Solve the function.

OpenStudy (anonymous):

\[6(x+h)^2-\frac{(x+h)}{h}\rightarrow6x^2+12xh+h^2-\frac{x}{h}+1\]

OpenStudy (anonymous):

Cyter: I'm confused Blank: That is the wrong answer

OpenStudy (anonymous):

Rightly so! Made huge mistaje \[(6(x+h)^2-(x+h))-\frac{(6x^2-x)}{h}\rightarrow6h^2-\frac{6x^2}{h}+12xh+\frac{x}h-h+6x^2-x\]

OpenStudy (anonymous):

If this what you meant to write \[\frac{f(x+h)-f(x)}{h}\] then the solution would be; \[6h+12x-1\] if you take \[\lim_{h\to0}6h+12x-1=12x-1\]

OpenStudy (anonymous):

Which is the derivative of f(x) \[\frac{dy}{dx}=f'(x)=12x-1\]

OpenStudy (anonymous):

thanks for the help